OpenStudy (anonymous):

Will give you medals Solve for t. P = irt A. t =p/ir B. t =ir/p C. t = Pir D. t = P – i – r

OpenStudy (anonymous):

A.

OpenStudy (anonymous):

A P=irt Divide both sides of the equation by ir P/ir=irt/ir and the equation will be P/ir=t since on the right side the the ir in the numerator and the ir in the denominator will have cancelled each other. Remember, if you do it to one side of the equation, you must do it to the other side.

OpenStudy (anonymous):

You must divide both sides of the equation by ir

OpenStudy (anonymous):

Sapp

OpenStudy (anonymous):

Sapphire

OpenStudy (anonymous):

wasnt i talking to you yesterday